News by Date
News by Category
Solar System
White Dwarfs
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Galaxy Clusters
Cosmology/Deep Field
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
Web Shortcuts
Chandra Blog
RSS Feed
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal (opens in new window)
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Operations CXO Status Report

Saturday 28 August 1999 3.00am EST

The following is an overview of the events that occurred over the last 24 hours.

The P8 load continued to execute as expected. A segmented maneuver scheduled for 10:28-11:33pm on 8/27 completed successfully and confirmed confirmed that the corrective actions taken after the safe mode entry were effective. The maneuver was performed during a communications pass and with appropriate engineering staff on console. The maneuver took Chandra to the bright star Capella in preparation for the deployment of the High Energy Transmission Grating (HETG).

The HETG was opened successfully at 1:35 am on 8/28. The procedure opened both launch locks and commanded the grating to the inserted position via hardware commands. All engineering data appeared nominal during the grating insertion. Data received by ACIS-S within 15 minutes showed the anticipated "X" pattern from the HEG and MEG facets and indicated that the HETG was inserted and operating as expected. A short time later features were evident in the spectrum. The cheer from the science room was audible in the control room! The grating will remain in place for 40 hours to obtain calibration data with ACIS-S then be retracted.

Return to Status Reports